# Optimal Error Analysis of Channel Estimation for IRS-assisted MIMO Systems

### v1

@@ Appendix A Introduction of the TT decomposition
where tensor factors ${\bm{B}}_{d}\in\mathbb{C}^{r_{d-1}\times N_{d}\times r_{d}},d=1,\dots,D$ with $r_{0}=r_{D}=1$. Thus, the TT format can be represented by $D$ tensor factors $\{{\bm{B}}_{d}\}_{d\geq 1}$, with a total of $O(D\overline{N}\overline{r}^{2})$ parameters, where $\overline{N}=\max_{d}N_{d}$ and $\overline{r}=\max_{d}r_{d}$. In addition, for any two TT format tensors ~,^$\widetilde{\mathcal{B}},\widehat{\mathcal{B}}\in\mathbb{C}^{N_{1}\times\cdots\times N_{D}}$ with factors $\{\widetilde{\bm{B}}_{d}(s_{d})\in\mathbb{C}^{\widetilde{r}_{d-1}\times\widetilde{r}_{d}}\}$ and $\{\widehat{\bm{B}}_{d}(s_{d})\in\mathbb{C}^{\widehat{r}_{d-1}\times\widehat{r}_{d}}\}$, each element of the summation $\mathcal{B}=\widetilde{\mathcal{B}}+\widehat{\mathcal{B}}$ can be represented by
$\mathcal{B}(s_{1},\dots,s_{D})=\begin{bmatrix}\widetilde{\bm{B}}_{1}(s_{1})\!\!\!\!&\widehat{\bm{B}}_{1}(s_{1})\end{bmatrix}\begin{bmatrix}\widetilde{\bm{B}}_{2}(s_{2})\!\!\!\!&{\bm{0}}\\ {\bm{0}}\!\!\!\!&\widehat{\bm{B}}_{2}(s_{2})\end{bmatrix}$
$\cdots\begin{bmatrix}\widetilde{\bm{B}}_{D-1}(s_{D-1})\!\!\!\!&{\bm{0}}\\ {\bm{0}}\!\!\!\!&\widehat{\bm{B}}_{D-1}(s_{D-1})\end{bmatrix}\begin{bmatrix}\widetilde{\bm{B}}_{D}(s_{D})\\ \widehat{\bm{B}}_{D}(s_{D})\end{bmatrix},$ (30)
@@
Since any tensor can be decomposed in the TT format (29) with sufficiently large TT ranks [37, Theorem 2.1], the decomposition of a tensor $\mathcal{B}$ into the form (29) is generally not unique: not only are the factors ${\bm{B}_{d}(:,s_{d},:)}$ not unique, but also the dimensions of these factors can vary. To introduce the factorization with the smallest possible dimensions $\bm{r}=(r_{1},\ldots,r_{D-1})$, for convenience, for each $d$, we put $\{{\bm{B}_{d}(:,s_{d},:)}\}_{s_{d}=1}^{N_{d}}$ together into the following two forms
$L(\bm{B}_{d})=\begin{bmatrix}\bm{B}_{d}(1)\\ \vdots\\ \bm{B}_{d}(N_{d})\end{bmatrix}\in\mathbb{C}^{(r_{d-1}N_{d})\times r_{d}},$ (31)
$R(\bm{B}_{d})=\begin{bmatrix}\bm{B}_{d}(1)&\cdots&\bm{B}_{d}(N_{d})\end{bmatrix}\in\mathbb{C}^{r_{d-1}\times(N_{d}r_{d})},$ (32)
@@
Appendix B Proof of Theorem 1
Proof.
By substituting the concentration inequality from [64, Theorem 2.3] with the concentration inequality for complex-valued subgaussian random variables, when (9) holds true, we can assert with a probability of $1-e^{-cT}$ that:
[…]
Due to $\|\mathcal{X}(\mathcal{B})\|_{F}^{2}=\sum_{p,k}\|\mathcal{X}(\mathcal{B})(p,k,:)\|_{2}^{2}$ and $\|\mathcal{B}\|_{F}^{2}=\sum_{p,k}\|\mathcal{B}(p,k,:)\|_{F}^{2}$, this completes the proof. ∎
@@ Appendix C Proof of Theorem 2
Using (7), we have
[…]
@@
[…]
@@
[…]
@@
According to (11), we can directly obtain
@@
[…]
where $\bm{X}^{*}$ is the conjugate matrix of $\bm{X}$ and $\mathcal{W}\times_{3}^{2}\bm{X}^{*}=\sum_{t=1}^{T}\mathcal{W}(:,:,t)\bm{X}^{*}(:,t)$. According to [48, eq.(44)], the last line follows $\|\mathcal{H}\|_{F}=\|\bm{H}_{2}\|_{F}\leq 1$ for a left-orthogonal TT form in Appendix A.
@@
Therefore, we can construct an $\epsilon$-net $\{\mathcal{H}^{(1)},\ldots,\mathcal{H}^{(n_{1}n_{2})}\}$ with covering number
$n_{1}n_{2}\leq\bigg{(}\frac{4+\epsilon}{\epsilon}\bigg{)}^{LPN+UMN}$ (40)
for any TT format tensor $\mathcal{H}=[\bm{H}_{1},\widetilde{\mathcal{S}},\bm{H}_{2}]\in\mathbb{B}_{2N,\widetilde{\mathcal{S}}}$.
@@
$[\widetilde{\bm{H}}_{1},\widetilde{\mathcal{S}},\widetilde{\bm{H}}_{2}]=\operatorname*{\text{arg~{}max}}_{\|\bm{H}_{1}\|\leq 1,\|\bm{H}_{2}\|_{F}\leq 1,\atop\|L(\widetilde{\mathcal{S}})\|\leq 1}1/T|\langle[\bm{H}_{1},\widetilde{\mathcal{S}},\bm{H}_{2}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|,$ (41)
$A:=\frac{1}{T}|\langle[\widetilde{\bm{H}}_{1},\widetilde{\mathcal{S}},\widetilde{\bm{H}}_{2}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|.$ (42)
Using $\mathcal{I}$ to denote the index set $[n_{1}]\times[n_{2}]$, then according to the construction of the $\epsilon$-net, there exists $p=(p_{1},p_{2})\in\mathcal{I}$ such that
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
where $c_{i},i=3,4,5$ are constants. Note that $P(F)$ follows Theorem 1.
@@
^$\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}\leq O\bigg{(}\frac{\gamma\sqrt{(1+\delta_{UM})(LPN+UMN)}}{(1-\delta_{UM})\sqrt{T}}\bigg{)}.$ (49)
Appendix D Proof of Theorem 3

### v2

@@ Appendix A Introduction of the TT decomposition
where tensor factors ${\bm{B}}_{d}\in\mathbb{C}^{r_{d-1}\times N_{d}\times r_{d}},d=1,\dots,D$ with $r_{0}=r_{D}=1$. Thus, the TT format can be represented by $D$ tensor factors $\{{\bm{B}}_{d}\}_{d\geq 1}$, with a total of $O(D\overline{N}\overline{r}^{2})$ parameters, where $\overline{N}=\max_{d}N_{d}$ and $\overline{r}=\max_{d}r_{d}$. In addition, for any two TT format tensors $\widetilde{\mathcal{B}},\widehat{\mathcal{B}}\in\mathbb{C}^{N_{1}\times\cdots\times N_{D}}$ with factors $\{\widetilde{\bm{B}}_{d}(s_{d})\in\mathbb{C}^{\widetilde{r}_{d-1}\times\widetilde{r}_{d}}\}$ and $\{\widehat{\bm{B}}_{d}(s_{d})\in\mathbb{C}^{\widehat{r}_{d-1}\times\widehat{r}_{d}}\}$, each element of the summation $\mathcal{B}=\widetilde{\mathcal{B}}$$\widehat{\mathcal{B}}$ can be represented by
$\hskip 0.0pt\mathcal{B}(s_{1},\dots,s_{D})$ $\!\!\!\!=\!\!\!\!$ $\begin{bmatrix}\widetilde{\bm{B}}_{1}(s_{1})\!\!\!\!&\widehat{\bm{B}}_{1}(s_{1})\end{bmatrix}\begin{bmatrix}\widetilde{\bm{B}}_{2}(s_{2})\!\!\!\!&{\bm{0}}\\ {\bm{0}}\!\!\!\!&\widehat{\bm{B}}_{2}(s_{2})\end{bmatrix}\cdots$ (31)
$\hskip-28.45274pt\begin{bmatrix}\widetilde{\bm{B}}_{D-1}(s_{D-1})\!\!\!\!&{\bm{0}}\\ {\bm{0}}\!\!\!\!&\widehat{\bm{B}}_{D-1}(s_{D-1})\end{bmatrix}\begin{bmatrix}\widetilde{\bm{B}}_{D}(s_{D})\\ \widehat{\bm{B}}_{D}(s_{D})\end{bmatrix}$,
@@
Since any tensor can be decomposed in the TT format (30) with sufficiently large TT ranks [38, Theorem 2.1], the decomposition of a tensor $\mathcal{B}$ into the form (30) is generally not unique: not only are the factors $\bm{B}_{d}(:,s_{d},:)$} not unique, but also the dimensions of these factors can vary. To introduce the factorization with the smallest possible dimensions $\bm{r}=(r_{1},\ldots,r_{D-1})$, for convenience, for each $d$, we put $\{{\bm{B}_{d}(:,s_{d},:)}\}_{s_{d}=1}^{N_{d}}$ together into the following two forms
$\hskip-28.45274ptL(\bm{B}_{d})=\begin{bmatrix}\bm{B}_{d}(1)\\ \vdots\\ \bm{B}_{d}(N_{d})\end{bmatrix}\in\mathbb{C}^{(r_{d-1}N_{d})\times r_{d}},$ (32)
$\hskip-28.45274ptR(\bm{B}_{d})=\begin{bmatrix}\bm{B}_{d}(1)&\cdots&\bm{B}_{d}(N_{d})\end{bmatrix}\in\mathbb{C}^{r_{d-1}\times(N_{d}r_{d})},$ (33)
@@
Appendix B Nonconvex Iterative Optimization Algorithms
In this section, we restate the nonconvex iterative algorithms used to estimate the channel matrices within the framework of the TT-based ToT regression model. For a single-hop IRS-assisted MIMO system, we directly adopt the bilinear alternating least squares (BALS) algorithm [31, Algorithm 2]. The computational complexity of each iteration of the BALS method, measured in terms of multiplication operations, is given by $O\big(TUMN+TLPKN+KTN\cdot\min\{KT,N\},LPKN\cdot\min\{LPK,N\}\big)$. In contrast, for a multi-hop IRS-assisted MIMO system, a closed-form solution for the alternating least squares method is not tractable. Therefore, we primarily consider the alternating gradient descent (AGD) method proposed in [77]. Note that we ignore the direct transmission path and focus on minimizing the following loss function:
$\!\!\!\!=\!\!\!\!$ $\frac{1}{T}\|\mathcal{X}([\widehat{\bm{B}}_{D},\mathcal{S}_{D},\dots,\widehat{\bm{B}}_{1},\mathcal{S}_{1},\widehat{\bm{B}}_{0}])-\widetilde{\mathcal{Y}}\|_{F}^{2}.$
where $\mu$ is a step size and the Wirtinger gradient $\nabla_{\bm{B}_{d}^{*}}g(\bm{B}_{D},\dots,\bm{B}_{0})$ is defined as:
$\hskip-14.22636pt\nabla_{\bm{B}_{d}^{*}}g(\bm{B}_{D},\dots,\bm{B}_{0})=\frac{1}{T}\sum_{k=1}^{K}\mathcal{S}_{d+1}^{\rm H}(:,k,:)\bm{B}_{d+1}^{\rm H}\cdots\mathcal{S}_{D}^{\rm H}(:,k,:)$
$\hskip 5.69046pt\cdot\bm{B}_{D}^{\rm H}\bigg(\mathcal{X}([\bm{B}_{D},\mathcal{S}_{D}(:,k,:),\dots,\bm{B}_{1},\mathcal{S}_{1}(:,k,:),\bm{B}_{0}])$
$\hskip 5.69046pt-\widetilde{\mathcal{Y}}(:,k,:)\bigg)\bm{X}^{\rm H}\bm{B}_{0}^{\rm H}\cdots\bm{B}_{d-1}^{\rm H}\mathcal{S}_{d}^{\rm H}(:,k,:).$ (35)
It is worth noting that AGD in this scenario operates without coefficient constraints, with the only requirement being $T=\Omega(UM)$ to ensure the RIP condition holds. Defining $N=\max_{d}N_{d}$, the overall computational complexity of each AGD iteration, measured in terms of multiplication operations, is $O\big(KD(DN^{3}+(N+T)LP+UMT(N+LP))\big)$.
@@ Appendix C Proof of Theorem 1
By substituting the concentration inequality from [65, Theorem 2.3] with the concentration inequality for complex-valued subgaussian random variables, when (9) in the main paper holds true, we can assert with a probability of $1-e^{-cT}$ that:
$(1-\delta_{UM})\|\mathcal{B}(p,k,:)\|_{F}^{2}$ $\!\!\!\!\leq\!\!\!\!$ $\frac{1}{T}\|\mathcal{X}(\mathcal{B})(p,k,:)\|_{2}^{2}$ (36)
$\!\!\!\!\leq\!\!\!\!$ $(1+\delta_{UM})\|\mathcal{B}(p,k,:)\|_{F}^{2}.$
Due to $\|\mathcal{X}(\mathcal{B})\|_{F}^{2}=\sum_{p,k}\|\mathcal{X}(\mathcal{B})(p,k,:)\|_{2}^{2}$ and $\|\mathcal{B}\|_{F}^{2}=\sum_{p,k}\|\mathcal{B}(p,k,:)\|_{F}^{2}$, this completes the proof for complex-valued subgaussian random variables.
In addition, it is straightforward to see that when $T\geq UM$, the normalized DFT matrix satisfies the RIP. ∎
Appendix D Proof of Theorem 2
Proof.
Using (7) in the main paper, we have
$0$ $\!\!\!\!\leq\!\!\!\!$ $\frac{1}{T}\|\mathcal{X}(\mathcal{B}^{\star})-\mathcal{Y}\|_{F}^{2}-\frac{1}{T}\|\mathcal{X}(\widehat{\mathcal{B}})-\mathcal{Y}\|_{F}^{2}$ (37)
$\!\!\!\!=\!\!\!\!$ $\frac{1}{T}\|\mathcal{X}(\mathcal{B}^{\star})-\mathcal{X}(\mathcal{B}^{\star})-\mathcal{W}\|_{F}^{2}$
@@ Appendix D Proof of Theorem 2
$\!\!\!\!=\!\!\!\!$ $\frac{2}{T}\operatorname{Re}\left\{\langle\mathcal{X}(\mathcal{B}^{\star})+\mathcal{W},\mathcal{X}(\widehat{\mathcal{B}}-\mathcal{B}^{\star})\rangle\right\}$
@@
$\!\!\!\!=\!\!\!\!$ $\frac{2}{T}\operatorname{Re}\left\{\langle\mathcal{W},\mathcal{X}(\widehat{\mathcal{B}}-\mathcal{B}^{\star})\rangle\right\}-\frac{1}{T}\|\mathcal{X}(\widehat{\mathcal{B}}-\mathcal{B}^{\star})\|_{F}^{2},$
@@
According to (11) in the main paper, we can directly obtain
@@
$\!\!\!\!=\!\!\!\!$ $\frac{2\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}}{T}\max_{\mathcal{H}\in\mathbb{B}_{2N,\mathcal{S}},\|\mathcal{H}\|_{F}\leq 1}|\langle\mathcal{H},\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|$
$\!\!\!\!=\!\!\!\!$ $\frac{2\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}}{T}\hskip-14.22636pt\max_{\|\bm{H}_{1}\|\leq 1,\|\bm{H}_{2}\|_{F}\leq 1,\atop\|L(\widetilde{\mathcal{S}})\|\leq 1}\hskip-21.33955pt|\langle[\bm{H}_{1},\widetilde{\mathcal{S}},\bm{H}_{2}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|,$
where $\bm{X}^{*}$ is the conjugate matrix of $\bm{X}$ and $\mathcal{W}\times_{3}^{2}\bm{X}^{*}=\sum_{t=1}^{T}\mathcal{W}(:,:,t)\bm{X}^{*}(:,t)$. According to [51, eq.(44)], the last line follows $\|\mathcal{H}\|_{F}=\|\bm{H}_{2}\|_{F}\leq 1$ for a left-orthogonal TT form in Appendix A of the main paper.
@@
Therefore, we can construct an $\epsilon$-net $\{\mathcal{H}^{(1)},\ldots,\mathcal{H}^{(n_{1}n_{2})}\}$ with covering number $n_{1}n_{2}\leq(\frac{4+\epsilon}{\epsilon})^{LPN+UMN}$ for any TT format tensor $\mathcal{H}=[\bm{H}_{1},\widetilde{\mathcal{S}},\bm{H}_{2}]\in\mathbb{B}_{2N,\widetilde{\mathcal{S}}}$.
@@
$\hskip-34.14322pt[\widetilde{\bm{H}}_{1},\widetilde{\mathcal{S}},\widetilde{\bm{H}}_{2}]=\hskip-14.22636pt\operatorname*{\text{arg~max}}_{\|\bm{H}_{1}\|\leq 1,\atop\|\bm{H}_{2}\|_{F}\leq 1,\|L(\widetilde{\mathcal{S}})\|\leq 1}\hskip-11.38092pt\frac{1}{T}|\langle[\bm{H}_{1},\widetilde{\mathcal{S}},\bm{H}_{2}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|,$ (43)
$\hskip-34.14322ptA:=\frac{1}{T}|\langle[\widetilde{\bm{H}}_{1},\widetilde{\mathcal{S}},\widetilde{\bm{H}}_{2}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|.$ (44)
Using $\mathcal{I}$ to denote the index set $n_{1}]\times[n_{2}]$, then according to the construction of the $\epsilon$-net, there exists $p=(p_{1},p_{2})\in\mathcal{I}$ such that
@@
$A$ $\!\!\!\!\leq\!\!\!\!$ $\frac{1}{T}|\langle[\bm{H}_{1}^{(p_{1})},\widetilde{\mathcal{S}},\bm{H}_{2}^{(p_{2})}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|$ (46)
@@
$\!\!\!\!=\!\!\!\!$ $\frac{1}{T}|\langle[\bm{H}_{1}^{(p_{1})},\widetilde{\mathcal{S}},\bm{H}_{2}^{(p_{2})}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|$
@@
$\!\!\!\!\leq\!\!\!\!$ $\frac{1}{T}|\langle[\bm{H}_{1}^{(p_{1})},\widetilde{\mathcal{S}},\bm{H}_{2}^{(p_{2})}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|+2\epsilon A$
$\!\!\!\!=\!\!\!\!$ $\frac{1}{T}|\langle[\bm{H}_{1}^{(p_{1})},\widetilde{\mathcal{S}},\bm{H}_{2}{(p_{2})}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|+\frac{A}{2}.$
@@
Furthermore, under the event $F:=\{\mathcal{X}\text{ satisfies $UM$-RIP}$  $\text{ with constant $\delta_{UM}$}\}$, which implies that $\frac{1}{T}\|\mathcal{X}(\mathcal{H}^{(p)})\|_{F}^{2}\leq(1+\delta_{UM})\|\mathcal{H}^{(p)}\|_{F}^{2}$. Plugging this together with the fact $\|\mathcal{H}^{(p)}\|_{F}\leq 1$ into the above further gives
@@
$\operatorname{\mathbb{P}}\left(A\geq t|F\right)$ $\!\!\!\!\leq\!\!\!\!$ $\operatorname{\mathbb{P}}\left(\max_{p_{1},p_{2}}\frac{1}{T}|\langle\mathcal{H}^{(p)},\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|\geq\frac{t}{2}|F\right)$ (49)
$\!\!\!\!\leq\!\!\!\!$ $\big(\frac{4+\epsilon}{\epsilon}\big)^{2LPN+2UMN}e^{-\frac{Tt^{2}}{8(1+\delta_{UM})\gamma^{2}}}$
$\!\!\!\!\leq\!\!\!\!$ $e^{-\frac{Tt^{2}}{8(1+\delta_{UM})\gamma^{2}}+c_{1}(LPN+UMN)},$
@@
$\!\!\!\!\geq\!\!\!\!$ $\operatorname{\mathbb{P}}\left(A\leq\frac{c_{2}\sqrt{(1+\delta_{UM})(LPN+UMN)}}{\sqrt{T}}\gamma\cap F\right)$
$\!\!\!\!\geq\!\!\!\!$ $P(F)\operatorname{\mathbb{P}}\left(A\leq\frac{c_{2}\sqrt{(1+\delta_{UM})(LPN+UMN)}}{\sqrt{T}}\gamma|F\right)$
$\!\!\!\!\geq\!\!\!\!$ $(1-e^{-c_{3}UM})(1-e^{-c_{4}(LPN+UMN)})$
$\!\!\!\!\geq\!\!\!\!$ $1-2e^{-c_{5}(LPN+UMN)},$
where $c_{i},i=3,4,5$ are constants. Note that $P(F)$ follows Theorem 1 in the main paper.
@@
$\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}=O\bigg(\frac{\gamma\sqrt{(1+\delta_{UM})(LPN+UMN)}}{(1-\delta_{UM})\sqrt{T}}\bigg).$ (51)
Appendix E Proof of (II-C) in the main paper
@@
Similar to the analysis presented in Appendix D, we focus on deriving an upper bound for the term $\frac{2}{T}|\langle\mathcal{W},\mathcal{X}(\widehat{\mathcal{B}}-\mathcal{B}^{\star})\rangle|$. To this end, we first rewrite the term as follows:
$\frac{2}{T}|\langle\mathcal{W},\mathcal{X}(\widehat{\mathcal{B}}-\mathcal{B}^{\star})\rangle|$ $\!\!\!\!=\!\!\!\!$ $2/T\hskip-22.76228pt\max_{\mathcal{B}\in\{\mathcal{C}:\mathcal{C}\in\mathbb{C}^{LP\times K\times UM},\atop\|\mathcal{C}\|_{F}\leq\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}\}}\hskip-22.76228pt|\langle\mathcal{B},\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|.$ (52)
Without loss of generality, we assume $\widehat{\mathcal{B}},\mathcal{B}^{\star}\in\mathbb{B}=\{\mathcal{B}=[\bm{H}_{1}\bm{\Sigma}_{1}\bm{H}_{2},\mathcal{S},\bm{G}_{1}\bm{\Sigma}_{2}\bm{G}_{2}]:\bm{\Sigma}_{1}\in\mathbb{C}^{F_{1}\times F_{1}},\bm{\Sigma}_{2}\in\mathbb{C}^{F_{2}\times F_{2}},\|\bm{\Sigma}_{1}\|_{F}\leq L_{1},\|\bm{\Sigma}_{2}\|_{F}\leq L_{2},\|\bm{\Sigma}_{1}\|_{0}=s_{1},\|\bm{\Sigma}_{2}\|_{0}=s_{2}\ \text{are unknown and}\ \mathcal{S}\in\mathbb{C}^{N\times K\times N},\bm{H}_{1}\in\mathbb{C}^{LP\times F_{1}},\bm{H}_{2}\in\mathbb{C}^{F_{1}\times N},\bm{G}_{1}\in\mathbb{C}^{N\times F_{2}}\ \text{and}\ \bm{G}_{2}\in\mathbb{C}^{F_{2}\times UM}\ \text{are known}\}$. According to [98, eq.(2)], we can construct an $\epsilon_{1}$-net $\{\bm{\Sigma}_{1}^{(1)},\dots,\bm{\Sigma}_{1}^{(n_{1})}\}$ with the covering number $n_{1}\leq(\frac{CF_{1}L_{1}}{s_{1}\epsilon_{1}})^{s_{1}}$ where $C$ is a positive constant, for the set of factors $\{\bm{\Sigma}_{1}\in\mathbb{C}^{F_{1}\times F_{1}}:\|\bm{\Sigma}_{1}\|_{F}\leq L_{1}\ \text{and}\ \|\bm{\Sigma}_{1}\|_{0}=s_{1}\}$ such that
$\sup_{\bm{\Sigma}_{1}\in\mathbb{C}^{F_{1}\times F_{1}}:\|\bm{\Sigma}_{1}\|_{F}\leq L_{1}\atop\ \text{and}\ \|\bm{\Sigma}_{1}\|_{0}=s_{1}}\min_{p_{1}\leq n_{1}}\|\bm{\Sigma}_{1}-\bm{\Sigma}_{1}^{(p_{1})}\|\leq\epsilon_{1}.$ (53)
Similarly, we can construct an $\epsilon_{2}$-net $\{\bm{\Sigma}_{2}^{(1)},\dots,\bm{\Sigma}_{2}^{(n_{2})}\}$ with the covering number $n_{2}\leq(\frac{CF_{2}L_{2}}{s_{2}\epsilon_{2}})^{s_{2}}$ for the set of factors $\{\bm{\Sigma}_{2}\in\mathbb{C}^{F_{2}\times F_{2}}:\|\bm{\Sigma}_{2}\|_{F}\leq L_{2}\ \text{and}\ \|\bm{\Sigma}_{2}\|_{0}=s_{2}\}$ such that
$\sup_{\bm{\Sigma}_{2}\in\mathbb{C}^{F_{2}\times F_{2}}:\|\bm{\Sigma}_{2}\|_{F}\leq L_{2}\atop\ \text{and}\ \|\bm{\Sigma}_{2}\|_{0}=s_{2}}\min_{p_{2}\leq n_{2}}\|\bm{\Sigma}_{2}-\bm{\Sigma}_{2}^{(p_{2})}\|\leq\epsilon_{2}.$ (54)
Therefore, we can construct an $\epsilon$-net $\{\mathcal{H}^{(1)},\ldots,\mathcal{H}^{(n_{1}n_{2})}\}$ with covering number $n_{1}n_{2}\!\!\leq\!\!(\frac{CF_{1}L_{1}}{s_{1}\epsilon_{1}})^{s_{1}}(\frac{CF_{2}L_{2}}{s_{2}\epsilon_{2}})^{s_{2}}$ for any TT format tensor $\mathcal{H}=[\bm{H}_{1}\bm{\Sigma}_{1}\bm{H}_{2},\mathcal{S},\bm{G}_{1}\bm{\Sigma}_{2}\bm{G}_{2}]$.
Denote by $B$ the value of (52), i.e.,
$\hskip-22.76228pt[\bm{H}_{1}\widehat{\bm{\Sigma}}_{1}\bm{H}_{2},\mathcal{S},\bm{G}_{1}\widehat{\bm{\Sigma}}_{2}\bm{G}_{2}]=\hskip-14.22636pt\operatorname*{\text{arg~max}}_{\mathcal{B}\in\{\mathcal{C}:\mathcal{C}\in\mathbb{C}^{LP\times K\times UM},\atop\|\mathcal{C}\|_{F}\leq\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}\}}\frac{1}{T}|\langle\mathcal{B},\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|,$
$\hskip-22.76228ptB:=\frac{1}{T}|\langle[\bm{H}_{1}\widehat{\bm{\Sigma}}_{1}\bm{H}_{2},\mathcal{S},\bm{G}_{1}\widehat{\bm{\Sigma}}_{2}\bm{G}_{2}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|.$
Using $\mathcal{I}$ to denote the index set $n_{1}]\times[n_{2}]$, then according to the construction of the $\epsilon$-net, there exists $p=(p_{1},p_{2})\in\mathcal{I}$ such that
$\|\widehat{\bm{\Sigma}}_{1}-\bm{\Sigma}_{1}\|\leq\epsilon_{1},\ \ \ \text{and}\ \ \ \|\widehat{\bm{\Sigma}}_{2}-\bm{\Sigma}_{2}\|\leq\epsilon_{2}.$ (55)
Then taking $\epsilon_{1}=\frac{L_{1}}{4}$ and $\epsilon_{2}=\frac{L_{2}}{4}$ gives
$B$ $\!\!\!\!\leq\!\!\!\!$ $\frac{1}{T}|\langle[\bm{H}_{1}\bm{\Sigma}_{1}^{(p_{1})}\bm{H}_{2},\mathcal{S},\bm{G}_{1}\bm{\Sigma}_{2}^{(p_{2})}\bm{G}_{2}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|$
$+\frac{1}{T}|\langle[\bm{H}_{1}\widehat{\bm{\Sigma}}_{1}\bm{H}_{2},\mathcal{S},\bm{G}_{1}\widehat{\bm{\Sigma}}_{2}\bm{G}_{2}]$
$-[\bm{H}_{1}\bm{\Sigma}_{1}^{(p_{1})}\bm{H}_{2},\mathcal{S},\bm{G}_{1}\bm{\Sigma}_{2}^{(p_{2})}\bm{G}_{2}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|$
$\!\!\!\!\leq\!\!\!\!$ $\frac{1}{T}|\langle[\bm{H}_{1}\bm{\Sigma}_{1}^{(p_{1})}\bm{H}_{2},\mathcal{S},\bm{G}_{1}\bm{\Sigma}_{2}^{(p_{2})}\bm{G}_{2}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|$
$+\frac{\epsilon_{1}}{T}|\langle[\bm{H}_{1}\frac{(\widehat{\bm{\Sigma}}_{1}-\bm{\Sigma}_{1}^{(p_{1})})}{\|\widehat{\bm{\Sigma}}_{1}-\bm{\Sigma}_{1}^{(p_{1})}\|_{F}}\bm{H}_{2},\mathcal{S},\bm{G}_{1}\widehat{\bm{\Sigma}}_{2}\bm{G}_{2}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|$
$+\frac{\epsilon_{2}}{T}|\langle[\bm{H}_{1}\bm{\Sigma}_{1}^{(p_{1})}\bm{H}_{2},\mathcal{S},\bm{G}_{1}\frac{(\widehat{\bm{\Sigma}}_{2}-\bm{\Sigma}_{2}^{(p_{2})})}{\|\widehat{\bm{\Sigma}}_{2}-\bm{\Sigma}_{2}^{(p_{2})}\|_{F}}\bm{G}_{2}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|$
$\!\!\!\!=\!\!\!\!$ $\frac{1}{T}|\langle[\bm{H}_{1}\bm{\Sigma}_{1}^{(p_{1})}\bm{H}_{2},\mathcal{S},\bm{G}_{1}\bm{\Sigma}_{2}^{(p_{2})}\bm{G}_{2}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|+\frac{B}{2}.$
We further assume that $\bm{H}_{1}\bm{\Sigma}_{1}^{(p_{1})}\bm{H}_{2},\mathcal{S},\bm{G}_{1}\bm{\Sigma}_{2}^{(p_{2})}\bm{G}_{2}]\in\mathbb{B}\cap\{\mathcal{B}=[\bm{H}_{1}\bm{\Sigma}_{1}\bm{H}_{2},\mathcal{S},\bm{G}_{1}\bm{\Sigma}_{2}\bm{G}_{2}]:\|\mathcal{B}\|_{F}\leq\|\widehat{\mathcal{B}}-\mathcal{B}\|_{F}\}$. Following the same analysis of (51), with high probability, we have
$\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}=O\bigg(\frac{\gamma\sqrt{(1+\delta_{UM})(s_{1}\log(\frac{F_{1}}{s_{1}})+s_{2}\log(\frac{F_{2}}{s_{2}}))}}{(1-\delta_{UM})\sqrt{T}}\bigg).$
Appendix F Proof of Theorem 3
Proof.